OpenStudy (anonymous):

X+Y=-2 NEED 3 POINTS FOR GRAPH

OpenStudy (anonymous):

Set it up as: y = -2 - x Now using x = 0, 1, and 2 we get: y = -2 - 0 = -2 y = -2 - 1 = -3 y = -2 - 2 = -4 So our three points are (0, -2), (1, -3), and (2, -4)

OpenStudy (anonymous):

HOW DO I DO -2Y=X=8

OpenStudy (anonymous):

-2Y=X+8

OpenStudy (anonymous):

start by dividing everything by -2 so the new equation is: y = -x/2 - 4

OpenStudy (anonymous):

x = 0 , y = -4 x = 1 , y = -4.5 x = 2 , y = -5

OpenStudy (anonymous):

THANKS